OpenStudy (anonymous):

in this polynomial x^2+?x-25 what numbers replaces the ? to produce a differenceof squares can someone check my answer

OpenStudy (anonymous):

would it it be 10

OpenStudy (anonymous):

10x

OpenStudy (mertsj):

Note that it is not a difference so it cannot be the DIFFERENCE of two squares unless ? is some negative number.

OpenStudy (anonymous):

so its wrong

OpenStudy (mertsj):

Suppose ? = -1. what would the problem be then?

OpenStudy (anonymous):

i made a mistake so now my answer would be right

OpenStudy (mertsj):

I hope so.

OpenStudy (mathstudent55):

Is the second x supposed to have an exponent of 25? If so, there is no number that can replace the ? to make this a difference of squares, since x^25 is not a square.

OpenStudy (anonymous):

the problem is x2+?x-25

OpenStudy (mathstudent55):

Then it's easy. The answer is 0.

OpenStudy (anonymous):

wtf ahahhahaah i see

OpenStudy (mathstudent55):

To have a difference of squares you need a subtraction involving only 2 squares. You can't have a middle term. If the ? is 0, then you have \(x^2+~?x-25 \) \(= x^2 + 0x - 25\) \(= x^2 - 25\) and Presto! you have a difference of squares.